if z to the power 11 + z bar = 0 then no. of solutions is ?

Grade:11

1 Answers

Vikas TU
3 years ago
Let z=x+iy
now according to question
z^2+zbar=0
(x+iy)^2+(x-iy)=0
x^2+2ixy-y^2+x-iy=0
(x^2+x-y^+i(2xy-y)=0
now
2xy-y=0 and (x^2+x-y^2)=0
2xy-y=0
here x=1/2
put this (x^2+x-y^2)=0
y^2=3/4
y=+_roo3/2
hence (1/2, root3/2)
and (1/2,-root3/2)